# CYCle: Choosing Your Collaborators Wisely to Enhance Collaborative Fairness in Decentralized Learning

### v1

@@ C.4 Experiments on CIFAR-100
We perform the same experiment outlined in Table 1 and Figure 2 on CIFAR-100 dataset. Figure 10 and Table 6 depicts the performance of each participant within the $N=5$ group across the various data partitioning strategies detailed in Section 5.1. Similar to the observations in Figure 2, in the imbalanced (0.8, 1) scenario, the FedAvg algorithm results in a decline in accuracy for $\mathcal{P}_{1}$, indicative of a negative collaboration gain. Following this, we compile the mean validation accuracy (MVA), mean collaboration gain (MCG), and collaboration gain spread (CGS) for each method and data splitting scenario in Table 4.
@@
Partition Number SA FedAvg VPDL CYCle
Homogeneous 1 $48.85$ $70.25$ $57.05$ $56.10$
2 $47.65$ $70.10$ $56.00$ $57.05$
3 $48.45$ $69.55$ $57.30$ $57.10$
4 $50.25$ $69.85$ $57.25$ $58.55$
5 $50.60$ $70.35$ $56.10$ $59.00$
Dirichlet (0.5) 1 $39.40$ $69.20$ $48.15$ $50.35$
2 $31.30$ $69.25$ $46.80$ $44.15$
3 $38.00$ $67.80$ $47.00$ $49.80$
4 $39.50$ $67.70$ $50.10$ $52.00$
5 $31.05$ $67.60$ $39.75$ $43.60$
Partition Number SA FedAvg VPDL CYCle
Imbalanced (0.8, 1) 1 $74.50$ $71.80$ $70.05$ $75.30$
2 $20.80$ $73.45$ $27.40$ $29.85$
3 $21.25$ $72.10$ $28.40$ $29.70$
4 $22.25$ $72.10$ $28.50$ $30.20$
5 $20.35$ $73.55$ $29.70$ $30.40$
Imbalanced (0.35, 2) 1 $61.45$ $70.00$ $62.10$ $67.15$
2 $59.30$ $70.60$ $59.35$ $65.95$
3 $29.65$ $70.55$ $40.15$ $39.80$
4 $28.10$ $70.95$ $41.05$ $40.30$
5 $32.05$ $69.55$ $39.75$ $41.95$
Partition Number SA FedAvg VPDL CYCle
Imbalanced (0.6, 1) 1 $68.25$ $70.80$ $65.80$ $70.80$
2 $29.90$ $70.90$ $39.30$ $39.00$
3 $30.15$ $71.80$ $38.70$ $40.70$
4 $28.50$ $71.35$ $39.45$ $38.05$
5 $29.40$ $69.15$ $39.80$ $39.85$
C.5 Imbalanced Data Study
In this study, we explore a range of values for the parameters $\kappa$ and $m$ in imbalanced settings, as detailed in Section 5.1. The purpose of this variation is to demonstrate the effectiveness of our proposed CYCle method in ensuring positive collaboration gains for each participant, showcasing a higher degree of fairness in comparison to FedAvg and VPDL. For a comprehensive view of these results, please see Figure 11 and Table 5.
@@ C.5 Imbalanced Data Study
Standalone ($\mathcal{B}_{n}$) $83.42$ $83.67$ $83.15$ $83.13$ $83.63$ $83.34$
Flip rate = 0.0 $86.12$ $87.03$ $86.33$ $85.62$ $86.10$ $86.28$
Flip rate = 0.2 $84.95$ $84.5$ $84.5$ $86.45$ $60.95$ $85.10$
Flip rate = 0.5 $84.25$ $84.65$ $84.65$ $83.25$ $55.80$ $84.20$
Flip rate = 1.0 $83.9$ $84.5$ $83.95$ $85.1$ $10.75$ $84.36$
C.6 Free Rider Study
Table 8: Validation accuracies of participants ($\mathcal{P}_{1}$, …, $\mathcal{P}_{5}$) with $\mathcal{P}_{4}$ and $\mathcal{P}_{5}$ having varying rates of label flipping. The last column refers to the average accuracy of honest participants ($\mathcal{P}_{1}$, …, $\mathcal{P}_{3}$).
@@ C.6 Free Rider Study
Table 7, aligned with the scenario depicted in Figure 5, details the values of $\mathcal{B}_{n}$ and $\mathcal{A}_{n}$. It reveals that, in the presence of flipped labels, the collaboration gain declines, reflecting a loss from a participant that holds 20% of the data.
@@
Figure 12: Heatmap visualization of reputation scores when the labels at $\mathcal{P}_{4}$ and $\mathcal{P}_{5}$ get corrupted by label flipping.
In Figure 12 and Table 8, we examine the outcomes using our CYCle approach under the varying degrees of label flipping (ranging from 0.0 to 1.0) when two participants ($\mathcal{P}_{4}$ and $\mathcal{P}_{5}$) are free riders. The findings show that the honest participants ($\mathcal{P}_{1}$, $\mathcal{P}_{2}$, and $\mathcal{P}_{3}$) maintain positive collaboration gains, despite the presence of malicious users / free riders. It’s observed that when the flip rate is greater than 0, the collaboration gain decreases relative to scenarios where the flip rate is 0. This decline is linked to the fact that these three participants collectively hold only 60% of the total dataset. In this experiment, we use CIFAR-10 dataset and a homogeneous data splitting approach, with other parameters remaining consistent with those detailed in Section 5.1.
C.7 Trade-off between MCG and CGS
Figure 13 presents a plot depicting the MCG and CGS values for 10 different data-split scenarios, which are specified in Table 1 and Table 5. It must be emphasized that the outcomes of our CYCle approach fall predominantly towards the bottom-right corner of the plot, indicating higher MCG and lower CGS values. On average, CYCle achieves the highest MCG and lowest CGS across the 10 scenarios. While competing methods (CFFL, RFFL, and CGSV) may slightly perform better than CYCle in one metric, they perform significantly worse in the other.
@@ C.8 Hyperparameter Sensitivity
C.9 Additional Results
The accuracy metrics for the individual participants for the experiments on CIFAR-10 dataset (corresponding to Figures 2 and 3) are summarized in Table 9 for greater clarity. These results clearly demonstrate the negative gain experienced by participant $\mathcal{P}_{1}$ in the imbalanced settings under most existing FL algorithms and how the proposed CYCle approach overcomes this problem.
@@ C.9 Additional Results
Partition No./Metric SA FedAvg VPDL CYCle
Homogeneous 1 $83.42$ $91.02$ $85.82$ $86.12$
@@
Partition No./Metric SA CFFL RFFL CGSV CYCle
Homogeneous 1 $60.98$ $62.05$ $61.65$ $63.32$ $72.42$

### v2

@@ G.4 Experiments on CIFAR-100
We perform the same experiment outlined in Table 5.1 and Figure 3 on the CIFAR-100 dataset. Our goal is to assess the effectiveness and fairness of CYCle under increased label granularity and distributional heterogeneity. Figure 17 and Table 8 depict the performance of each participant within the $N=5$ group across the various data partitioning strategies detailed in Section 5.1. Similar to the observations in Figure 3, in the imbalanced (0.8, 1) scenario, the FedAvg algorithm results in a decline in accuracy for $\mathcal{P}_{1}$, indicative of a negative collaboration gain. Following this, we compile the mean validation accuracy (MVA), mean collaboration gain (MCG), and collaboration gain spread (CGS) for each method and data splitting scenario in Table 9.
As in CIFAR-10 experiments, we observe that FedAvg often achieves the highest mean validation accuracy (MVA), particularly in homogeneous and moderately in imbalanced settings. However, this comes at a significant fairness cost. VPDL, while privacy-preserving and decentralized, underperforms in both MVA and MCG across all scenarios. Although it avoids extreme gain inequality (moderate CGS), it fails to leverage meaningful collaboration in data-heterogeneous environments, largely due to its uniform treatment of client contributions.
@@
G.5 Imbalanced Data Study
In this study, we explore a range of values for the parameters $\kappa$ and $m$ in imbalanced settings, involving five participants ($N=5$), as detailed in Section 5.1. The purpose of this variation is to demonstrate the effectiveness of our proposed CYCle method in ensuring positive collaboration gains for each participant, showcasing a higher degree of fairness in comparison to FedAvg and VPDL. For a comprehensive view of these results, please see Figure 18 and Table 10. Table 10 reports performance across five different imbalance settings using three metrics: MVA, MCG, and CGS. FedAvg suffers from a higher CGS and, in some scenarios, a large variance in individual participant outcomes, indicating fairness issues. VPDL offers stronger privacy guarantees but significantly underperforms in both accuracy and fairness. CYCle, in contrast, achieves a compelling balance. It consistently improves fairness over both baselines, as evidenced by its lower CGS across all five settings. This suggests that CYCle more equitably distributes collaborative benefits.
These findings confirm that CYCle not only preserves collaboration benefits but also effectively mitigates the challenges of unequal data distributions, an essential feature for practical decentralized learning algorithms.
G.6 Free Rider Study
In Table 19 and Figure 19, we examine the outcomes using our CYCle approach under the varying degrees of label flipping (ranging from 0.0 to 1.0) when two participants ($\mathcal{P}_{4}$ and $\mathcal{P}_{5}$) are free riders. The findings show that the honest participants ($\mathcal{P}_{1}$, $\mathcal{P}_{2}$, and $\mathcal{P}_{3}$) maintain positive collaboration gains, despite the presence of malicious users / free riders. It’s observed that when the flip rate is greater than 0, the collaboration gain decreases relative to scenarios where the flip rate is 0. This decline is linked to the fact that these three participants collectively hold only 60% of the total dataset. In this experiment, we use the CIFAR-10 dataset and a homogeneous data splitting approach, with other parameters remaining consistent with those detailed in Section 5.1.
Table 11: Validation accuracies of participants ($\mathcal{P}_{1}-\mathcal{P}_{5}$) with $\mathcal{P}_{4}$ and $\mathcal{P}_{5}$ having varying rates of label flipping. The last column refers to the average accuracy of honest participants ($\mathcal{P}_{1}-\mathcal{P}_{3}$).
@@ G.6 Free Rider Study
Figure 19: Heatmap visualization of reputation scores with corrupted labels at $\mathcal{P}_{4}$ and $\mathcal{P}_{5}$.
G.7 Scalability to Larger Client Populations
To evaluate the applicability of CYCle beyond cross-silo federated learning, we conducted additional experiments on CIFAR-10 and CIFAR-100 with $N=20$ clients. Data was partitioned using a Dirichlet distribution with $\alpha=0.5$ to simulate non-i.i.d. heterogeneity across clients. We compare three methods: standalone (no collaboration), Gossip-SGD (with full connectivity), and CYCle.
(a) CIFAR-10
(b) CIFAR-100
Figure 20: Per-participant performance radar plots for $N=20$ clients on (a) CIFAR-10 and (b) CIFAR-100 datasets using three methods: standalone (no collaboration), Gossip-SGD (with a complete graph), and CYCle.
Figure 20 displays the per-participant accuracies on CIFAR-10 and CIFAR-100 datasets. In particular, clients $2,7,$ and $15$ exhibit significantly lower standalone performance (Figure 20(a)) – indicating limited data quality or quantity. CYCle is able to identify such low-contributing clients through a reputation mechanism and adjusts collaboration accordingly. This ensures that while these clients benefit from collaboration, they do not disproportionately influence others, thus maintaining fairness and robustness.
These results demonstrate that CYCle maintains robust fairness and stable performance when scaled to larger client populations beyond the cross-silo setting. In contrast to Gossip-SGD, which exhibits a degradation of fairness (indicated by a negative correlation in Table 12), CYCle achieves a significantly more equitable distribution of collaborative gains.
For CIFAR-10, both methods achieve similar MCG values (28.59 for Gossip-SGD vs. 27.87 for CYCle), indicating comparable average collaborative benefits. However, the collaboration gain spread (CGS) is considerably lower for CYCle ($6.06$ vs. $7.70$), showing that it reduces the inequality in the results of collaboration. Moreover, the strong positive correlation ($\rho=0.9353$) under CYCle implies that participants with more useful or valuable data tend to receive proportionally higher rewards, signaling a fairness-aligned reward mechanism. In contrast, Gossip-SGD shows a negative correlation ($\rho=-0.4637$), which implies misalignment between client contribution and received gains.
Dataset Method MCG CGS Pearson corr. $(\rho)$
CIFAR-10 Gossip-SGD 28.59 7.70 -0.4637
CYCle 27.87 6.06 0.9353
CIFAR-100 Gossip-SGD 22.68 2.12 -0.7297
CYCle 22.88 1.25 0.8164
Table 12: Comparison of collaborative fairness metrics for $N=20$ clients.
A similar trend holds for CIFAR-100. The MCG values are again comparable, but CYCle achieves a markedly lower CGS, indicating greater consistency in how benefits are distributed. Pearson’s correlation improves dramatically from $\rho=-0.7297$ (Gossip-SGD) to $\rho=0.8164$ (CYCle), further supporting that CYCle not only promotes fairness, but aligns collaboration rewards more effectively with actual client contribution.
We also note that CYCle can be naturally extended to scenarios with significantly larger client populations and partial participation. In such cases, a subset of clients can be sampled in each round and CYCle can be applied by constructing a subgraph over the active clients. This involves computing a sub-mixing matrix using only the participating clients and running the CYCle protocol locally on this induced set of clients. Reputation and gradient alignment remain local and round-specific, requiring no global state. This design makes CYCle compatible with the standard partial participation strategies commonly used in cross-device FL. However, since our focus is on cross-silo settings with full participation, we leave empirical evaluation under partial participation for future work.
In summary, CYCle preserves average performance while significantly improving fairness and reward alignment between participants. These gains are particularly important in federated settings with data heterogeneity and varying client quality, as they help prevent both under- and over-compensation in collaborative learning.
G.8 Trade-off between MCG and CGS
Figure 21 presents a plot depicting the MCG and CGS values for 10 different data-split scenarios, which are specified in Table 5.1 and Table 10. Each point in the plot corresponds to the average performance of a method under one scenario, with the $x$-axis representing the mean collaboration gain (MCG) and the $y$-axis representing the collaboration gain spread (CGS). The bottom right region of the plot, indicating high MCG and low CGS, represents the ideal operational regime, where all participants benefit meaningfully and fairly from collaboration. It must be emphasized that the results of our CYCle approach fall predominantly towards the bottom-right corner of the plot. On average, CYCle achieves the highest MCG and the lowest CGS in the 10 scenarios. While baseline methods such as CFFL, RFFL, and CGSV may outperform CYCle in one of the two metrics under certain scenarios, they exhibit a clear trade-off: improvements in collaborative gain often come at the cost of high inequality (larger CGS) and vice versa. For example, CGSV may occasionally achieve a higher MCG, but its CGS is substantially worse, indicating that a few participants benefit disproportionately, while others gain little or even suffer negative outcomes. In contrast, methods such as RFFL may better control CGS but fail to deliver sufficient gain, as reflected in lower MCG values.
The directionality of the arrows overlaid on the figure further emphasizes the optimization goal: to move toward the bottom right.
@@ G.9 Hyperparameter Sensitivity
These parameters play a well-defined and interpretable role in our framework through the mapping function (Equation 7). The parameters $\tau_{opt}$ and $\tau_{max}$ define a soft acceptance region over the similarity space (e.g., give full credit to highly aligned clients $h(s)=1$, while blocking collaboration with poorly aligned ones $h(s)=0$). As shown in Figure 22, we explored a range of reasonable values and found that going below $\tau_{opt}=0.25$ results in degraded collaboration fairness and utility, as unqualified updates are included too permissively. The selected values are therefore informed by both the empirical similarity statistics and the theoretical geometry of the collaboration space. For these reasons, we did not conduct exhaustive tuning and instead used principled, task-informed choices.
CYCle (Gossip-SGD).
The only hyperparameter in the gossip-based variant of our protocol is $\beta$, which controls the sharpness of the softmax function used to derive the communication probabilities from gradient alignment scores. Higher $\beta$ values assign greater weight to stronger alignments, enforcing more selective and fairness-oriented peer interactions, while lower values yield more uniform communication, resembling a fully connected topology. Figure 23(a) shows the performance across participants under varying $\beta\in\{5,15,35,50,100\}$. Although participant-level performance varies slightly, all configurations exhibit reasonable fairness, validating the robustness of CYCle under different settings. Figure 23(b) further illustrates the trade-off between utility (MCG) and fairness (Pearson correlation of pre- and post-collaboration accuracies). We observe that intermediate values of $\beta$ (e.g., 35-50) strike a favorable balance, delivering high collaboration gain without sacrificing fairness.
(a) Performance of CYCle under different $\beta$ parameters $\in\{5,15,35,50,100\}$.
(b) MCG vs. correlation coefficient.
Figure 23: Study on the impact of parameter $\beta$ on performance and analysis of the utility versus fairness trade-off using the CYCle algorithm.
G.10 Additional Results
The accuracy metrics for the individual participants for the experiments on the CIFAR-10 dataset (corresponding to Figures 3 and 4) are summarized in Table 13 for greater clarity. These results clearly demonstrate the negative gain experienced by the participant $\mathcal{P}_{1}$ in the imbalanced settings under most existing FL algorithms and how the proposed CYCle approach overcomes this problem.
@@ G.10 Additional Results
Split No./Metric SA FedAvg VPDL CYCle
Homogeneous
1 $83.42$ $91.02$ $85.82$ $86.12$
@@
Split No./Metric SA CFFL RFFL CGSV CYCle
Homogeneous
1 $60.98$ $62.05$ $61.65$ $63.32$ $72.42$